# Evolutionary Prediction Games

### v1

@@ Abstract
When users decide whether to use a system based on the quality of predictions they receive, learning has the capacity to shape the population of users it serves—for better or worse. This work aims to study the long-term implications of this process through the lens of evolutionary game theory. We introduce and study evolutionary prediction games, designed to capture the role of learning as a driver of natural selection between groups of users, and hence a determinant of evolutionary outcomes. Our main theoretical results show that: (i) in settings with unlimited data and compute, learning tends to reinforce the survival of the fittest, and (ii) in more realistic settings, opportunities for coexistence emerge. We analyze these opportunities in terms of their stability and feasibility, present several mechanisms that can sustain their existence, and empirically demonstrate our findings using real and synthetic data.
Machine Learning, ICML
@@ 1 Introduction
Accurate predictions have potential to improve many aspects of our lives, whether as individuals, as groups, or as a society. From everyday content recommendation to life-changing medical diagnoses, higher quality predictions in human-facing tasks often lead to better decisions and therefore better outcomes for users. A reasonable conclusion is that we should work hard to further push the envelope on predictive accuracy across applications. Considerable efforts in machine learning research and practice are devoted precisely to this purpose. But at the same time, there is a growing recognition that blindly promoting accuracy in social contexts can have unexpected, and sometimes undesired, consequences. It is therefore important to consider not only how much, but also how, predictive accuracy is attained.
Concerns regarding the application of machine learning in society typically consider its possible ill effects on a given population of users. Taking this notion one step further, here we argue that learning can have the potency to determine what this population is—or will be. Our main observation is that when the benefit to users depends on the accuracy of the predictions they receive, then consequently, so will the willingness of those users to use the system.Since users’ decisions of whether or not to use a system are in aggregate and over time precisely what forms the population of system users, learning gains influence over the population’s eventual composition. Our goal is to study the capacity of learning to do so—a point we believe is often overlooked, but carries significant implications.
Our key observation is that not all users can be equally happy; or more technically, that in typical settings, not all groups of users in a population will get equally accurate predictions. If we make the plausible assumption that groups who get higher accuracy are more likely to grow (e.g., by satisfied users inviting their peers), whereas groups who get lower accuracy are prone to shrink (e.g., by users leaving or convincing others to leave too), then over time, learning will come to shape the relative proportions of groups in the population. From a learning perspective, once the population changes, it becomes beneficial to retrain the model on new data from the updated distribution. This creates a feedback loop in which changes to the model induce changes in the population, which then lead to further changes to the model, ad infinitum. We are interested in understanding the long-term outcomes of this process, both in terms of the attainable accuracy and the resulting composition of the population.
@@
Towards this, we adopt a novel evolutionary perspective and model the joint dynamics of learning and user choices through the lens of natural selection. Under the assertion that some degree of predictive error always exists, our key modeling point is that accuracy becomes, in effect, a scarce resource over which different groups in the population ‘compete’. By associating each group’s accuracy with its evolutionary fitness, we obtain an evolutionary process in which learning is the driver of selective pressure, and consequently, a determinant of long-term population outcomes.
Our main innovation is to model this process using a population game, which we adopt from the field of evolutionary game theory (e.g., Hofbauer & Sigmund, 1998), which studies the dynamics of populations of agents driven by myopic interaction rules. Population games are a core component of this framework, defining the evolutionary fitness of each group as a function of the overall population state. Formalizing evolution as a game connects key properties of natural selection processes (e.g. stationary points) with properties of the corresponding population game (e.g. Nash equilibria). Such links have previously been established for a large family of games and dynamics (see, e.g., Sandholm, 2010), and successfully used for explaining a wide array of evolutionary phenomena from survival of the fittest and kin selection to altruistic behavior and mating rituals.
Population games are useful for studying the limiting behavior of global population dynamics driven by local interactions, such as imitation, word-of-mouth, or social learning. This makes them well-suited for adaptation to our setting. Building on this as motivation, we propose a new form of population game, which we call an evolutionary prediction game, that models the coevolution of a (re)trained model and the population of users over which it makes predictions. Evolutionary prediction game are unique in that the interaction between the different groups are induced by solutions to a non-trivial optimization problem, namely loss minimization. This complexity introduces inherent challenges to characterizing the game’s outcomes, but at the same time, provides useful structure, which we exploit in our analysis.
A primary question that evolutionary game theory aims to answer is: which species survive? Typically, there are two possible outcomes of interest: survival of some species (and extinction of others), or coexistence. Evolutionary prediction games allow us to explore which of these outcomes materialize, and when, for a population of social groups in settings where learning acts as a selective force. In nature, most interactions result in the ‘survival of the fittest’, an outcome which is so pronounced that it is often associated with the mechanism of natural selection itself. Our first result establishes that this occurs in learning as well, in two settings: (i) when the model is trained only once at the onset (i.e., is not updated over time), and (ii) under the ideal conditions of infinite data and unlimited compute. While intuitive, proving this for evolutionary prediction games is technically challenging due to the complexity introduced by the loss minimization operator. Our main tool here is to show that evolutionary prediction games in such settings adhere to the structure of more general potential games (Sandholm, 2001). This allows us to connect evolutionary outcomes to extrema of scalar-valued functions over the simplex, providing interpretation and simplifying analysis.
One interesting point that our analysis reveals is that, in the idealized setting, although survival of the fittest is the likely outcome, coexistence becomes a possibility under retraining. Technically, this manifests as the (possible) existence of a mixed equilibrium, but which is unstable (and so will not materialize under plausible dynamics). Our next set of results then show that learning under realistic conditions, i.e., using finite data and/or limited compute, can stabilize such mixed equilibria, meaning that coexistence becomes possible (or even probable). We demonstrate several mechanisms through which this phenomenon can arise, all common in machine learning practice, namely: (i) the use of proxy losses (e.g., hinge loss or cross entropy), (ii) the need to overcome overfitting, and (iii) asymmetric label noise. We conduct experiments using synthetic and real data: the former to provide insight into the inner workings of these mechanisms, and the latter to demonstrate their capacity for coexistence under more realistic conditions and at scale.
Figure 1: Interaction model in a two-group setting. Population is a mixture ${\bm{p}}$ of groups; Classifier $h$ is learned using data from the mixture distribution $D_{\bm{p}}$; Evolutionary fitness is associated with prediction accuracy; Differences in fitness drive change in mixture coefficients; Possible long-term outcomes are dominance (only $A$ survives), extinction (only $B$ survives), or coexistence (both survive).
@@ 1.1 Related Work
Our work relates to the emerging literature on performative prediction (Perdomo et al., 2020), which studies learning in settings where model deployments affect the underlying data distribution, with emphasis on (re)training dynamics and equilibrium. A central effort in this area is to identify global properties (e.g., appropriate notions of smoothness) that guarantee convergence. Our work complements this effort by proposing structure, in the form of evolutionary forces. This allows us to study a rich class of stateful performative dynamics, which are notoriously more challenging (Brown et al., 2022), thanks to an efficient low-dimensional representation of the distribution map (see Section A.1).
Evolutionary game theory was originally developed to model biological natural selection (Maynard Smith & Price, 1973), but has also been used in sociology (Bisin & Verdier, 2001) and financial markets (Scholl et al., 2021).The evolutionary prediction games we introduce are novel since they are defined as implicit solutions to statistical optimization problems. Our results for optimal classifiers reflect the principle of competitive exclusion (Hardin, 1960), while our analysis of non-optimal classifiers relates to long-studied questions of coexistence (e.g., Chesson, 2000).
Our formulation of evolutionary prediction games gives rise to a notion of equilibrium that satisfies the fairness criterion of overall accuracy equality in expectation (e.g., Verma & Rubin, 2018). Thus, rather than encouraging fair outcomes, we study a setting in which fairness is emergent. This adds another dimension to the question of fair outcomes, since it asks not only whether equity holds for the observed groups, but also which groups will remain to be observed. It also suggests an intriguing counterfactual perspective: if we observe fair outcomes at present, could this be because some groups were historically driven out of the game? This complements other voiced concerns regarding long-term outcomes of fair learning (Hashimoto et al., 2018; Liu et al., 2018; D’Amour et al., 2020).
@@ 2 Setting
The core of our setting is based on a standard supervised learning setup in which examples describe user data. Denote features by $x\in\mathcal{X}$ and labels $y\in\mathcal{Y}$. For a given distribution $D$ over pairs $(x,y)$, and given a training set $S={\left\{{(x_{i},y_{i})}\right\}}_{i=1}^{n}\sim D^{n}$, the goal in learning is to use $S$ to find a predictor $h:\mathcal{X}\to\mathcal{Y}$ from a class $\mathcal{H}$ which minimizes the expected error under some loss function $\ell:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$. For concreteness, we focus primarily on classification tasks with the 0-1 loss, $\ell(\hat{y},y)=\mathds{1}\left\{{\hat{y}\neq y}\right\}$, whose minimization is equivalent to maximizing expected accuracy:
@@
Denote the learning algorithm by $\mathcal{A}(S)$, and by $h\sim\mathcal{A}(D)$ the process mapping distributions to learned classifiers, with randomness due to sampling. Since solving Eq. (1) is both computationally and statistically hard, practical learning algorithms often resolve to optimizing an empirical surrogate loss (e.g., hinge or cross-entropy). One question we will ask is how such compromises affect long-term outcomes.
@@
As data points $(x,y)$ are generated by users, the data distribution $D$ represents a population. We assume the population is a mixture of $K$ groups. Each group $k\in[K]$ is associated with a group-specific distribution $D_{k}$ over $(x,y)$ pairs, which remains fixed, and with a relative size $p_{k}$, which evolves over time. The overall data distribution is a mixture denoted by $D=D_{\bm{p}}=\sum_{k}p_{k}D_{k}$, where $p_{k}$ describes the relative proportion of group $k$ (such that $\sum_{k}p_{k}=1$ and $p_{k}\geq 0\,\,\forall k$). The population state vector ${\bm{p}}=(p_{1},\dots,p_{K})$ will be our main object of interest. Formally, the space of population states is the $(K-1)$-dimensional simplex $\Delta^{K}$. As ${\bm{p}}$ determines $D_{\bm{p}}$, for brevity we denote $h\sim\mathcal{A}({\bm{p}})=\mathcal{A}(D_{\bm{p}})$ and $\mathrm{acc}_{\bm{p}}(h)=\mathrm{acc}_{D_{\bm{p}}}(h)$. Our goal is to understand how changes in ${\bm{p}}$ affect learning, and in turn, how learning and its outcomes shape states ${\bm{p}}$.
@@
For a classifier $h\sim\mathcal{A}({\bm{p}})$, denote the marginal expected accuracy of $h$ on group $k$ by $\mathrm{acc}_{k}(h)=\mathrm{acc}_{D_{k}}(h)$. Informally, our central assumption on dynamics is that the proportion $p_{k}$ of each group $k$ grows if $\mathrm{acc}_{k}(h)$ is larger than others (on average), and shrinks if it is smaller. This aims to capture settings in which users benefit from accurate predictions, and so their tendency to use the system (or not) will depend on how well the system performs on other users in their cohort. Since $h$ is trained on the entire population, it is likely that some groups will benefit more (and therefore grow), and others less (and so will shrink). Thus, each deployment of a classifier induces change in ${\bm{p}}$, resulting in a form of distribution shift known as subpopulation shift (Yang et al., 2023). This, in turn, will influence the choice of classifier in the next time step—which forms a feedback loop between ${\bm{p}}$ and $h$. We propose to formalize and study such dynamics through the lens of evolutionary game theory, as described next.
@@ 3 Evolutionary Prediction Games
Figure 2: Graphical representation of two-group evolutionary prediction games (Definition 3.1). Groups are denoted by ${\left\{{A,B}\right\}}$. Solid lines describe expected group accuracies for each population state ${\bm{p}}$, dotted line describes population accuracy $\mathrm{acc}_{\bm{p}}(h)$, dots indicate Nash equilibria, and x-axis arrows indicate evolutionary stability. Each schematic diagram represents a possible scenario: (Left) State-dependent dominance. There are two stable equilibria, each dominated by a single group, with locally maximal population accuracy. Additionally, there is an unstable coexistence equilibrium. (Center Left) Stable beneficial coexistence. There is a stable coexistence equilibrium, with maximal population accuracy. (Center Right) Unstable beneficial coexistence. Coexistence has maximal population accuracy, but it is unstable. (Right) Total group dominance. There is a single equilibrium dominated by a single group.
Evolutionary Game Theory is a general framework for the analysis of natural selection processes (Maynard Smith & Price, 1973). The framework is built around population games – symmetric normal-form games for large number of agents, which model the interactions between $K$ species in a large mixed population. The population is modeled as a distribution over species ${\bm{p}}\in\Delta^{K}$. Each species $k\in[K]$ corresponds to a pure strategy in the population game, and is associated with a scalar fitness function, $F_{k}({\bm{p}})$, which quantifies its evolutionary fitness given population state ${\bm{p}}$ [^1] For example, in a biological context, evolutionary fitness $F_{k}({\bm{p}})$ may represent the average number of offspring who manage to reach maturity for an individual of species $k$. The fitness functions $F=\left(F_{1},\ldots,F_{K}\right)$ jointly define the population game, which then governs the dynamics of ${\bm{p}}$.
Learning as an evolutionary force.
We formalize the interplay between learning and user population dynamics through a novel type of population game, which we call an evolutionary prediction game. This game associates the fitness of each group with its expected marginal accuracy:
Definition 3.1 (Evolutionary Prediction Game).
@@
Together, the fitness functions $F=(F_{1},\dots,F_{K})$ define an evolutionary prediction game over the population.
The game captures an implicit interaction between groups: While the fitness of each group $F_{k}({\bm{p}})$ is given by its own accuracy, the classifier $h$ is trained on the entire population, and thus different groups interact through the learning process. An illustration of this idea is outlined in Figure 1.
Dynamics revisited.
Our analysis considers the joint evolution of population states and classifiers. As population games may induce dynamics in several ways, we maintain generality by making only two structural assumptions on dynamics, namely that: (i) changes in $p_{k}$ are positively correlated with fitness $F_{k}({\bm{p}})$, and (ii) Nash equilibria of the population game are rest points of the dynamics. These hold for many well-known families of dynamical processes, including replicator dynamics (Taylor & Jonker, 1978), Moran processes (Moran, 1958), and multiplicative weight updates (e.g., Kleinberg et al., 2009) – See definitions in Section B.1. For the learning process, we first consider a setting where the classifier is trained only once at the onset (Section 4.1), and then proceed to study retraining dynamics in which $h$ is retrained at each step to fit the new distribution $D_{\bm{p}}$ (Sections 4.25). This draws connections to the recent literature on performative prediction (Perdomo et al., 2020).
Evolutionary outcomes.
Denote the initial population state by ${\bm{p}}^{0}$, and the initial classifier by $h^{0}\sim\mathcal{A}({\bm{p}}^{0})$. We will be interested in characterizing the joint evolutionary outcomes of ${\bm{p}}$ and $h$. For the dynamics we consider, this takes the form of equilibrium points, denoted $({\bm{p}}^{*},h^{*})$. Games may have multiple equilibria, and the eventual outcome will depend on the group distributions, the learning algorithm, and the initial conditions. Our taxonomy of possible equilibria concerns three key properties:
  • Survival: For the population, some groups may become extinct (e.g., $p^{*}_{k}\to 0$ for some $k\in[K]$), while others may reach coexistence (e.g., $p^{*}_{k}\to 1/K$ for all $k\in[K]$).
  • Accuracy: For the classifier, natural selection may result in its performance improving over time, (i.e. $\mathrm{acc}(h^{*})>\mathrm{acc}(h^{0})$), deteriorating, or remaining unchanged.
  • Stability: For the system, some equilibrium points will be robust to small perturbation of ${\bm{p}}^{*}$ and remain stable, while for others this may cause dynamics to diverge.
When $K=2$, the probability simplex $\Delta^{K}$ is equivalent to the unit interval $[0,1]$, and evolutionary prediction games admit a convenient graphical representation, illustrating the relations between the different elements (see Figure 2).
Nash equilibria and fairness.
For Nash equilibria, we follow the definition given by Sandholm (2010). Let $F({\bm{p}})$ be a population game. For a given population state ${\bm{p}}$, denote $\operatorname*{support}({\bm{p}})={\left\{{k\mid p_{k}>0}\right\}}$. A state ${\bm{p}}^{*}$ is a Nash equilibrium of $F({\bm{p}})$ if it satisfies the following condition:

### v2

@@ Abstract
When a prediction algorithm serves a collection of users, disparities in prediction quality are likely to emerge. If users respond to accurate predictions by increasing engagement, inviting friends, or adopting trends, repeated learning creates a feedback loop that shapes both the model and the population of its users. In this work, we introduce evolutionary prediction games, a framework grounded in evolutionary game theory which models such feedback loops as natural-selection processes among groups of users. Our theoretical analysis reveals a gap between idealized and real-world learning settings: In idealized settings with unlimited data and computational power, repeated learning creates competition and promotes competitive exclusion across a broad class of behavioral dynamics. However, under realistic constraints such as finite data, limited compute, or risk of overfitting, we show that stable coexistence and mutualistic symbiosis between groups becomes possible. We analyze these possibilities in terms of their stability and feasibility, present mechanisms that can sustain their existence, and empirically demonstrate our findings.
@@ 1 Introduction
Accurate predictions have become essential for any platform that supports user decision-making. Improvements in prediction accuracy often directly translate to better quality of service with benefits to both the platform and its users. This is one reason why modern platforms ranging from content recommendation and online marketplaces to personalized education and medical services have come to rely on machine learning as their backbone, and are investing much effort and resources in continually improving their predictions. However, while generally beneficial, promoting accuracy blindly can have unexpected, and in some cases undesired, consequences. It is therefore important to understand the possible downstream and long-term effects of learning on social outcomes.
Conventional learning approaches aim to maximize accuracy on a given, predetermined data distribution. But in social settings, the distributions are composed of those users who choose to use the platform. When such choices depend on the quality of predictions, learning becomes a driver of user self-selection, and thus gains influence over its user population. This creates a feedback loop: model deployment shifts the population, and population changes trigger model retraining. We are interested in understanding the general tendencies and possible long-term outcomes of this process.
In particular, our focus is on user feedback dynamics in which accurate predictions encourage engagement or adoption, and prediction errors discourage them. Such dynamics arise across domains: accurate recommendations drive network-effect growth as users invite peers; successful content strategies on media platforms are mimicked by others; precise credit-risk models reduce premiums and therefore attract more clients to profitable loan programs; and hospitals with higher diagnostic accuracy draw more patients. Common to the above is that individual choices adhere to some form of group structure in the population, and that in aggregate, these groups tend to grow when prediction quality is higher. Note this notion of “group” is in itself quite flexible, as groups can represent different demographics, behaviors, or roles, and with memberships being either inherent or chosen.
@@
Focusing on user choices as the primary driving force, we seek to analyze the impact of learning on group proportions over time. For this purpose, we adopt a novel evolutionary perspective and model the joint dynamics of learning and user choices through the lens of natural selection. Under the assertion that some degree of predictive error always exists, our key modeling point is that accuracy becomes, in effect, a scarce resource over which different groups in the population ‘compete’. By associating each group’s accuracy with its evolutionary fitness, we obtain an evolutionary process in which learning is the driver of selective pressure, and consequently, a determinant of long-term population outcomes. We can then ask questions regarding the long-term tendencies of the population composition, overall and per-group accuracies, and the affect of different modeling and algorithmic choices on temporal trends and long-term evolutionary outcomes.
Contributions.
Our main conceptual innovation is to analyze these feedback loops using a population game—a core component of evolutionary game theory useful for studying the dynamics of agents driven by local interaction rules [33, 21, 43]. Evolutionary game theory relates between game-theoretic properties of population games (e.g., Nash equilibria), and the dynamics of a wide variety of natural selection dynamics, such as imitation, word-of-mouth influence, or social learning, all of which are supported by our analysis. This makes the framework well-suited for our setting, as it allows us to address the common thread in the wide variety of feedback loops described above in a unified way.
Towards this, we first propose a novel game, called an evolutionary prediction game, in which evolutionary fitness is associated with prediction accuracy. This allows us to study the co-evolution of a (re)trained model and the population of its users. We then analyze the structure of games induced by different learning algorithms and their corresponding dynamics, and characterize the possible outcomes under different settings. Our results include conditions under which ‘survival of the fittest’ is a likely outcome; mechanisms that nonetheless enable coexistence; a discussion of the role and effects of retraining; and connections between survival, accuracy, and (evolutionary) stability. We also discuss connections to fairness and social conservation.
We conclude by complementing our analysis with experiments using both synthetic and real data and coupled with simulated dynamics. Our empirical results shed light on when and how certain user groups are likely to dominate, disappear, or coexist, and demonstrate how different design choices can shape social outcomes—even if inadvertently. Together, these highlight the importance of understanding, anticipating, and accounting for the long-term effects of learning on its user population.
@@ 1.1 Related Work
Evolutionary game theory.
The use of game theory for studying evolutionary processes was originally developed to model biological natural selection [34, 33, 21, 47], but has also been used in domains such as economics [43, 14], social sciences [2, 31], and online learning dynamics [13, 5]. The games we propose are unique in this space as they are defined as implicit solutions to statistical optimization problems. Our results for oracle classifiers reflect the principle of competitive exclusion [17], while our analysis of non-optimal classifiers relates to long-studied questions of coexistence [e.g., 7].
Performativity.
Our work relates to the emerging field of performative prediction [39], which studies learning in settings where model deployments affect the underlying data distribution, with emphasis on (re)training dynamics and equilibrium. A central effort in this area is to identify general global properties (e.g., appropriate notions of smoothness) that guarantee convergence. Our work complements this effort by proposing structure, in the form of user self-selection [55, 22], which provides an efficient low-dimensional representation of the distribution map. This allows us to work in the notoriously challenging stateful performative setting [3], and to characterize stronger notions of stability.
Fairness.
The notion of equilibrium in evolutionary prediction games relates to the fairness criterion of overall accuracy equality [51]. However, rather than proposing a way to encourage fair outcomes, we study a setting in which fairness is emergent. This adds another dimension to the question of fairness, since it asks not only whether equity holds for the observed groups, but also which groups will remain to be observed. It also suggests an intriguing counterfactual perspective: if we observe fair outcomes at present, could this be because some groups were historically driven out of the game? This complements other voiced concerns regarding long-term outcomes of fair learning [19, 32, 10, 40].
Additional discussion of related work is provided in Appendix B.
Figure 1: Natural selection in a two-group setting. Population is a mixture ${\bm{p}}$ of groups; Classifier $h$ is learned using data from the mixture distribution $D_{\bm{p}}$; Evolutionary fitness is associated with prediction accuracy; Differences in fitness drive change in mixture coefficients; Possible long-term tendencies are dominance (only $A$ survives), extinction (only $B$ survives), or coexistence (both survive).
@@ 2 Setting
The core of our setting is based on a standard supervised learning setup in which examples describe user data. Denote features by $x\in\mathcal{X}$ and labels $y\in\mathcal{Y}$. For a given distribution $D$ over pairs $(x,y)$, and given a training set $S={\left\{{(x_{i},y_{i})}\right\}}_{i=1}^{n}\sim D^{n}$, the goal in learning is to use $S$ to find a predictor $h:\mathcal{X}\to\mathcal{Y}$ from a class $\mathcal{H}$ which minimizes the expected error under some loss function $\ell:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$. For concreteness, we focus on classification tasks with the 0-1 loss, $\ell(\hat{y},y)=\mathds{1}\left\{{\hat{y}\neq y}\right\}$, whose minimization is equivalent to maximizing expected accuracy:
@@
Denote the learning algorithm by $\mathcal{A}(S)$, and by $h\sim\mathcal{A}(D)$ the process mapping distributions to learned classifiers, with randomness due to sampling. Since solving Eq. (1) is both computationally and statistically hard, practical learning algorithms often resolve to optimizing an empirical surrogate on the sampled data. One question we will ask is how such compromises affect long-term outcomes.
@@
As data points $(x,y)$ are generated by users, the data distribution $D$ represents a population. We assume the population is a mixture of $K$ groups. Each group $k\in[K]$ is associated with a group-specific distribution $D_{k}$ over $(x,y)$ pairs, which remains fixed, and with a relative size $p_{k}$, which evolves over time. The overall data distribution is a mixture denoted by $D=D_{\bm{p}}=\sum_{k}p_{k}D_{k}$, where $p_{k}$ is the current relative proportion of group $k$ (such that $\sum_{k}p_{k}=1$ and $p_{k}\geq 0\,\,\forall k$). Given some $h$, the marginal expected accuracy on group $k$ is $\mathrm{acc}_{k}(h)=\mathrm{acc}_{D_{k}}(h)$. The population state vector ${\bm{p}}=(p_{1},\dots,p_{K})$ will be our main object of interest. As ${\bm{p}}$ encodes group proportions, the space of population states is the $(K-1)$-dimensional simplex $\Delta^{K}$, and since ${\bm{p}}$ determines $D_{\bm{p}}$, for brevity we denote $h\sim\mathcal{A}({\bm{p}})=\mathcal{A}(D_{\bm{p}})$ and $\mathrm{acc}_{\bm{p}}(h)=\mathrm{acc}_{D_{\bm{p}}}(h)$.
@@
When a classifier is deployed, users make decisions according to the quality of the predictions they receive. Collectively, individual responses reshape the population composition ${\bm{p}}$, resulting in a subpopulation shift [52]. The updated population then serves as the training distribution for the next classifier—which forms a feedback loop between ${\bm{p}}$ and $h$. For training, we consider a simple procedure of repeated training on the current distribution. In terms of user behavior, our framework accommodates a wide family of dynamics, including reproduction, network effects, imitation, social influence, rational decision-making, and competition, as alluded to in Section 1. In Appendix C we define these formally and provide additional examples. In Section 3 we make our assumptions explicit.
Outcomes.
Denote the initial population state by ${\bm{p}}^{0}$, and the initial classifier by $h^{0}\sim\mathcal{A}({\bm{p}}^{0})$. We consider dynamics which evolve towards fixed points, and denote the corresponding point by ${\bm{p}}^{*}$. Denote the classifier by $h^{*}\sim\mathcal{A}({\bm{p}}^{*})$. We characterize fixed points in terms of three key properties:
  • Population composition: For the population, dynamics may act as a force diminishing certain groups (i.e., driving towards a state satisfying $p^{*}_{k}=0$ for some $k\in[K]$). In such cases, we say that the population is dominated by the remaining groups. When the population composition at rest has at least two groups (i.e., $p^{*}_{k}>0$ for multiple values of $k$), we say that these groups coexist.
  • Accuracy: For the classifier, dynamics may result in overall performance improving over time, (i.e., $\mathrm{acc}(h^{*})>\mathrm{acc}(h^{0})$), deteriorating, or remaining unchanged.
  • Stability: For the joint system, some fixed points ${\bm{p}}^{*}$ are stable and attract all neighboring states, while other fixed points are unstable and may cause dynamics to diverge under small perturbations.
@@ 3 Evolutionary Prediction Games
Figure 2: Evolutionary dynamics for two groups, induced by an oracle linear classifier (Section 4). (Left) Data distributions in feature space. Dashed line demonstrates the optimal linear classifier $h_{\bm{p}}$ for the uniform mixture ${\bm{p}}=(0.5,0.5)$. (Center) Evolutionary prediction game $F_{k}({\bm{p}})=\mathrm{acc}_{k}(h_{\bm{p}})$. The game has two stable equilibria with single-group dominance, and one unstable coexistence equilibrium. Overall population accuracy $\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$ is convex, and maximized at the boundries. (Right) Replicator dynamics induced by the game, for various initial states ${\bm{p}}^{0}$. Populations evolve towards fixed points with single-group dominance.
To analyze the feedback loop, we leverage evolutionary game theory, which is a general framework for the analysis of natural selection processes [34]. The framework models evolutionary interactions between individuals from $K$ groups in a large mixed population: Given a population state ${\bm{p}}\in\Delta^{K}$, each group $k\in[K]$ is associated with a scalar fitness function, $F_{k}({\bm{p}})$, which quantifies the average evolutionary fitness of individuals in that group. Together, the fitness functions $F=\left(F_{1},\ldots,F_{K}\right)$ define a population game $F:\Delta^{K}\to\mathbb{R}^{K}$, which governs the dynamics of ${\bm{p}}$.
Formally, population games are symmetric normal-form games for a continuum of agents. Evolutionary game theory uses population games to map myopic, often local agent decisions to group-level dynamics governing the evolution of group proportions ${\bm{p}}$ over time. Modeling the game at the group level serves as useful abstraction for studying the general tendencies and possible long-term outcomes of local interactions in aggregate. Note the notion of "group" is quite flexible and can support memberships that are either inherent (e.g., by demographics) or chosen (e.g., service provider).
Accuracy as evolutionary fitness.
We formalize the interplay between learning and population dynamics through a novel type of population game—an evolutionary prediction game. The game associates the evolutionary fitness of each group with its expected marginal accuracy $\mathrm{acc}_{k}$ under a classifier $h$ trained on data sampled from the mixture $D_{\bm{p}}$:
Definition 1 (Evolutionary Prediction Game).
@@
Together, the tuple $F({\bm{p}})=(F_{1}({\bm{p}}),\dots,F_{K}({\bm{p}}))$ defines an evolutionary prediction game.
In the notation $F_{k}({\bm{p}})$, we assume for brevity that the learning algorithm $\mathcal{A}$ and group distributions $D_{k}$ are clear from context, and we state them explicitly otherwise.
Evolutionary prediction games capture an implicit interaction between groups: while each group’s fitness $F_{k}({\bm{p}})$ is given by the accuracy of its own members, the classifier $h_{\bm{p}}$ is trained using data from the entire population. Thus, the evolution of different groups becomes coupled through the learning process. For example, Figure 2 (Left) illustrates a binary classification setting over two groups ${\left\{{A,B}\right\}}$. Figure 2 (Center) shows the corresponding evolutionary prediction game $F({\bm{p}})$ induced by an oracle linear classifier (defined in Section 4). At ${\bm{p}}^{t}=(0.7,0.3)$, we have $F_{A}({\bm{p}}^{t})\approx 0.97$ and $F_{B}({\bm{p}}^{t})\approx 0.78$, and thus the fitness of group $A$ is higher when the classifier is trained on data sampled from $D_{{\bm{p}}^{t}}$.
Equilibrium and fairness.
Nash equilibrium in general population games is defined as follows:
Definition 2 (Nash equilibrium of a population game; e.g. [43]).
Let $F({\bm{p}})$ be a population game, and denote $\operatorname*{support}({\bm{p}})={\left\{{k\mid p_{k}>0}\right\}}$. A state ${\bm{p}}^{*}\in\Delta^{K}$ is a Nash equilibrium of $F({\bm{p}})$ if it satisfies: